The perimeter of a rectangular basketball court blocked off by students on a playground is 300 feet. If the length of the court is two feet more than triple the width, find the length and the width of the blacked off court.
Answer by checkley77(12569)

L=3W+2
2L+2W=300
2(3W+2)+2W=300
6W+4+2W=300
8W=300-4
8W=296
W=296/8
W=37 FT IS THE WIDTH.
L=3*37+2=113 FT FOR THE LENGTH.
PROOF.
2*37=2*113=300
74+226=300
300=300